Title

Article text.

INAF queue observing, 2016 December 5 - 6

Observer: R. Carini, A. Marchetti
Telescope Operator: G. Bechetti
Support Astronomer: O. Kuhn on call
Instrument: LUCI1, MODS1, MODS2

Summary

LUCI2 is not available

At sunset the sky is cloudy and there is west-southwest wind. For all the night there was clouds and strong wind until 3 a.m. (limiting the observation to targets in north-northeast direction), with horrible and strongly variable seeing, with peaks of 4". We experienced technical problems with the presets, both with LUCI and MODS: some scripts couldn't be sent, it is not clear if it is a telescope issue or a scripts issue, or something related to the LUCI2 to LUCI1 conversion.

Night duration (12\xB0 el twilight) 18:10-6:00 ( 01:45-12.45 UT) = 11h 50m

Observing time: 7h 56m (67%)

Weather time loss: 1h 37m (13%)

technical time loss 2h 17m (20%)

Data Summary

Project_RUN Objects Filter ExpTime Notes
Antoniucci PV Cep zj 300 sec seeing variable 1.7''-2.8''
Antoniucci V1118Ori zJ+HK   COMPLETED but seeing strongly variable
telluric Antoniucci zj+ Hk    
Annibali DD068 RunB dual grating 1 h 1 ob completed in MODS1 only.
Annibali Standard Feige34 dual grating    
CALIBRATIONS Annibali MODS 1 DD068      
CALIBRATIONS Antoniucci MODS1        

Details

01:04 the TO tries to collimate in the field of LUCI1 Piconcelli_27_S, with no success * 01:07 seeing on guider is 2.5"-3"

01:22 we move to the field of LUCI1 Antoniucci V512Per. The collimation had success *01:24 We send the script for the telluric. It is too cloudy, the telescope loses the guide star. The declination of the telluric is too low, the shell rips

01:25 resetting LUCI1 shell

01:27 we send the preset for the target of Antoniucci V512Per
  • 01:29 the telescope cannot find the guide star

  • 01:32 we send the preset again

  • 01:37 the telescope cannot find the guide star, but there's apparently no error in the script

  • 01:39 the SA is doublechecking the script, she finds no errors

  • 01:41 the seeing is 2.8"

  • 01:42 we try to re-select the script and change the guide star

  • 01:48 we send the preset again

  • 01:49 the telescope cannot find the guide star

01:53 we send the preset for Piconcelli_27_S. The telescope cannot find the guide star

01:54 we send the preset for the telluric of Piconcelli

01:58 there are probably some problems with LUCI1

02:00 the TO and SA begin tests to figure out the problem

02:46 we send again the preset for the Piconcelli telluric as a test. Now it works, but nothing was done or changed. Maybe it was due to weather conditions

02:52 we send the science of LUCI1 Antoniucci v512Per

  • 02:55 the telescope cannot find the guide star, because the pointing is off. It is a bad issue of the telescope that needs to be resolved, independent on the instrument

02:59 we send the science preset again. It fails. More tests

03:36 we change target and unbelievably LUCI works. Run PVCep for Antoniucci program
  • acquisition luci.20161206.0020-0023
  • seeing 2.85''
  • science luci.0024 (zj filter)
  • seeing variable 1.7''-2.8''
  • 03:56 guide star lost. the sky is covered.
  • tray again but it fails

4:02 the sky is covered, the TO closes the dome

05:39 the TO opens the dome

06:19 we slew to LUCI1 Antoniucci V1118Ori

06:23 the To collimates on the science field of LUCi1 Antoniucci V1118Ori

06:25 we send the science for LUCi1 Antoniucci V1118Ori
  • seeing less than 1" but the sky is thin.
  • acquisition luci1.20161206.0025-0028
  • science luci1.20161206.0029-0036

07:03 slew to telluric for Antoniucci V1118 ori
  • acquisition luci.0037-0042
  • seeing got worse to 2''-3''
  • the wind increases and comes from S-W
  • science luci.0043-0046

07:27 slew to Marziani A2
  • error : tilting the grating unit
  • run again the script
  • seeing about 3''.
  • error again. maybe the script is corrupted.

07:47 slew to telluric for the Piconcelli target A2 LUCI1 is stuck and the seeing is horrible, about > 3''

08:02 the TO suggest us to point at north to find a better seeing, at the point opposite of the wind
  • we try with telluric for the target A2 of Ricci
  • again LUCI1 stuck for the seeing: between 3''-4''

08:50 the seeing seems a little be better, the part of the sky were the seeing is better is E, or N-E, we switch to MODS

09:20 MODS is awake and open, but the TO sees no light

09:45 the problem has likely to do with the preset.

10:02 Problem is solved after calling the SA and sending the preset again. The seeing is bad, but we cannot send the back up programs, since they face the wind

10:04 We slew to Annibali program, which is visible and in good direction. Seeing is now about 1.5",1.6". the wind is about 10 m/s
  • we send the acquisition for Annibali DDO68 RunB
  • problems with the mask alignment in MODS2. We repeat the acquisition. We do not resolve the problem, and decide to send the science script anyway.
  • error: PARTNER INAF timed out after 60 seconds. We retry and everything goes fine. Seeing 1.4"-1.7", but sometimes peaks of 2.5". Probably a mass of warm air going up and interesting the telescope (there is a peak in temperature in the panel). Acquisition:

  • mods1r.20161206.0001-005.fits

  • mods2r.20161206.0001-005.fits

science: MODS2 is not taking any useful spectra, the mask is not aligned:

  • mods1b.20161206.0001-006.fits

  • mods1r.20161206.0006-0011.fits

  • mods2b.20161206.0001-006.fits BAD

  • mods2b.20161206.0006-0011.fits BAD

* 12:22* we acquire the standard for Annibali (Feige34), we try again the binocular mode
  • 11:18 seeing 1.4"
  • 11:20 seeing 1.2"
  • 11:29 seeing 1"
  • 11.36 seeing 0.9"-1"
  • 11:44 seeing less than 1.2", but sometimes it reaches 1.6
  • 12:07 seeing 0.7"-0.9"
  • 12:22 we send the script for the standard of Annibali (Feige34). We try again the binocular mode
  • 12:48 something went wrong with the acquisition in MODS2. It took about 40 images and it was not going to stop. We stopped the acquisition in MODS2.

  • mods2r.0007-0040.fits BAD

  • mods1r.0011-0018.fits (one extra image as a check after MODS2 went crazy)

12:52 we start the science of Feige34

  • mods1r.0019

  • mods1b.0007

13:30 we send the calibration script for Annibali DDO68. Problems with MODS2. We abort the script with MODS2. Error on MODS1. We send it again for MODS1 only. Error message again:

  • mods1r.0021-0025

  • mods1b.0009-0013

  • mods1r.0020 and mods1b.0008 are wrong

14:03 we send the calibrations for Antoniucci MODS1-2 LS slit 0.8" (flats and lamps).
  • problems with MODS2, many error messages. Apparently some problem with the script, that for some reason does not affect MODS1. Some flats and lamps were taken for MODS1.
  • we also take the imaging flats for MODS1 and MODS2, but MODS2 still gives error messages.

-- %USERSIG{AlidaMarchetti - 2016-12-04}%

Comments

 
Topic revision: r13 - 08 Dec 2016, RobertaCarini
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback